OpenStudy (anonymous):

Choose the equation of the line passing through the point (2, 6) and parallel to y = -3x - 4

OpenStudy (anonymous):

If the line is parallel, then it has the same slope. Put into point-slope form, the equation will be y-6=-3(x-2) or y=-3(x-2)+6.